Express the roots of the equation z3 −α3 = 0 in terms of α and w, where w is a complex cube root of unity. Use your answer to find the roots of the following equations in the form a + ib. a) z^2-27=0
1
Expert's answer
2020-05-20T18:54:41-0400
"z^3-\alpha^3=0"

"z^3-\alpha^3=(z-\alpha)(z^2+\alpha z+\alpha^2)"

Then


"(z-\alpha)(z^2+\alpha z+\alpha^2)=0"

"z^2+\alpha z+\alpha^2=0"


"z={-\alpha\pm\sqrt{\alpha^2-4\alpha^2}\over2}=\alpha({-1\pm i\sqrt{3}\over2})"

Cube Root of Unity Value


"w_1=1,\ real"

"w_2={-1- i\sqrt{3}\over2}, complex"


"w_3={-1+ i\sqrt{3}\over2}, \ complex"

"z_1=\alpha w_1=\alpha\cdot1=1\cdot\alpha+i\cdot0"

"z_2=\alpha w_2=\alpha\cdot{-1- i\sqrt{3}\over2}=-{1\over 2}\alpha-i\cdot{\alpha\sqrt{3}\over 2}"

"z_3=\alpha w_3=\alpha\cdot{-1+ i\sqrt{3}\over2}=-{1\over 2}\alpha+i\cdot{\alpha\sqrt{3}\over 2}"